OpenStudy (anonymous):

use the geometric mean to find the vaule of a12 a11=72 and a13=1/2

OpenStudy (anonymous):

a12 = root(a11 x a13)

OpenStudy (anonymous):

is the answer 143.49?

OpenStudy (anonymous):

whats the formula for geometric mean? i only know of sequence and progression

OpenStudy (anonymous):

6

OpenStudy (anonymous):

root (72 x 1/2)

OpenStudy (anonymous):

a12 = sqrt(a11 * a13) = sqrt(72*1/2) =sqrt(36) i.e 6 or -6